OpenStudy (anonymous):

Is the complex conjugate operator \(\hat{K}\) a Hermitian operator? As in does it satisfy this equation \[\int \Psi_1^* (\hat{K}\Psi_2) dx = int (\hat{K}\Psi_1)^* \Psi_2 dx\]

OpenStudy (anonymous):

i coud'nt understand any thing!!!!

OpenStudy (anonymous):

The complex conjugate operator takes the complex conjugate of soemthing like this \[\hat{K}(a+ib) = (a+ib)^*= a -ib\]